# High-resolution Giant Metrewave Radio Telescope Hi 21 cm imaging of the host galaxy of FRB 20250316A

### v1

@@ 3 Results and Discussion
Applying the above value and the G23 extinction model to our measured H line flux density, we obtain a dust-corrected H line luminosity of erg s-1. Using the calibration of R. C. Kennicutt & N. J. Evans (2012), this yields an SFR of yr-1 at the location of the eBOSS spectrum, kpc from the centre of the galaxy, for a Chabrier initial mass function. The eBOSS fibres are circular, with diameter , implying an SFR surface density of yr-1 kpc-2 for the region covered by the eBOSS fibre. This is an uncharacteristically high SFR surface density for the outskirts of a disk galaxy, and suggests that the recent burst of star-formation is likely to have been triggered by a merger or accretion event.
@@
We next used the “direct-Te” method to infer the metallicity at the location of the eBOSS spectrum (e.g. Y. I. Izotov et al., 2006; T. Jiang et al., 2019), using the extinction-corrected flux densities in the [Oiii]4363Å, [Oiii]4959Å, [Oiii]5007Å, and [Oii]3727Å lines. This yields . The metallicity at the galaxy centre is (L. J. Kewley et al., 2005). This would imply a metallicity gradient of dex kpc-1 towards the south-west. However, typical metallicity gradients in disk galaxies are dex kpc-1 (e.g. D. Zaritsky et al., 1994). A similar gradient, dex kpc-1, is obtained in NGC 4141 from the slightly sub-solar metallicity measured by Chime/Frb Collaboration (2025b) in the FRB environment. Therefore, the low metallicity at the location of the eBOSS spectrum instead suggests a recent addition of metal-poor gas, via either a merger or accretion from the circumgalactic medium. This fresh gas is likely to have resulted in the increased Hi column density seen in the south-western outskirts of the galaxy in Figs. 4[A] and [B], and to have triggered the recent star-formation activity, leading to the high inferred SFR surface density.
@@
A. K. Leroy et al. (2019) find that NGC 4141 lies dex above the local star-forming main sequence. Our estimate of the 1.38 GHz flux density ( mJy) allows us to obtain an independent radio-based estimate of the SFR of NGC 4141. Using the Tully-Fisher distance estimate of Mpc (R. B. Tully et al., 2023) for the galaxy, we obtain a radio-based SFR of year-1. This is lower than the SFR estimate of year-1 from the dust-corrected H line luminosity (L. J. Kewley et al., 2005). As noted by Chime/Frb Collaboration (2025b), this difference may arise due to a recent increase in the star-formation activity, as the radio SFR estimate is an average over 100 Myr, while the H-based estimate is an average over the last 10 Myr. The increased recent SFR may account for the detections of two Type-II SNe, an optical transient, an X-ray transient, and FRB 20250316A, all within the last two decades, as well as for the fact that the galaxy lies above the local main sequence.
@@
Overall, our GMRT Hi 21cm images find evidence for disturbances in the Hi distribution of NGC 4141 and the companion, LEDA 2582852, a nearby Hi cloud to the south-west without any discernible star-formation and at the velocities of the south-western side of the galaxy disk, and high Hi column densities in the south-western outskirts of the galaxy disk of NGC 4141. The SDSS spectrum of the south-western region of the galaxy yields a high SFR surface density and a low metallicity, while the recent ( Myr) SFR of NGC 4141 is higher than the radio-based SFR averaged over 100 Myrs. All of these indicate that NGC 4141 has recently acquired fresh gas, via either accretion or a merger, which compressed the Hi in the south-western part of the galaxy, triggering the starburst. This merger/accretion event may have also triggered additional star-formation activity in the disk of NGC 4141, resulting in the increased recent SFR and possibly in the formation of the stellar progenitor of FRB 20250316A, and the other transients. Earlier Hi 21cm mapping studies of FRB host galaxies have also yielded evidence for recent merger activity in a number of systems (e.g B. Kaur et al., 2022; M. Glowacki et al., 2024; H. Roxburgh et al., 2025). It is possible that the mergers triggered a burst of star-formation activity that resulted in the formation of the massive stellar progenitors of the ensuing radio bursts.
@@
Finally, our non-detection of continuum emission at the location of FRB 20250316A places a upper limit of 24 Jy on the 1.38 GHz flux density of any associated persistent radio source (PRS). This implies the upper limit of erg s-1 Hz-1 on the rest-frame 1.4 GHz specific luminosity of such a putative PRS. This is one of the strongest constraints on the radio luminosity of a PRS associated with a non-repeating FRB (e.g. C. J. Law et al., 2024b; Chime/Frb Collaboration, 2025b)
4 Summary
@@
We have used the GMRT Band-5 receivers to map Hi 21cm emission in and around NGC 4141, the host of FRB 20250316A, at , at angular resolutions of , i.e. spatial resolutions of kpc. The low redshift of NGC 4141 implies that this is the finest spatial resolution at which Hi 21cm emission has ever been mapped from an FRB host galaxy. The GMRT Hi 21cm images find evidence for a companion galaxy, LEDA 2582852, to the south-west of NGC 4141, a nearby Hi cloud also towards the south-west, high Hi column densities in the south-western outskirts of the galaxy disk, disturbances in the Hi distribution in both NGC 4141 and LEDA 2582852, and high velocity dispersions, km s-1, away from the central regions of NGC 4141. All of these suggest a recent merger event that compressed the Hi in the south-western disk, and resulted in a burst of star-formation activity in the galaxy. Consistent with this, an SDSS spectrum in the south-west yields evidence of a low metallicity, suggesting recent addition of metal-poor gas, via either accretion or a merger. The estimate of the recent SFR in NGC 4141, from the H line, is also higher than our estimate of the radio-based SFR, averaged over 100-Myr timescales, while the SDSS spectrum finds a high SFR surface density in the south-western disk. We conclude that a merger event within the last 100 Myr is likely to have compressed the Hi in the south-west of the galaxy and triggered the starburst activity. It is unclear whether the merger was with the Hi cloud identified kpc away from NGC 4141 or whether this cloud was ejected from NGC 4141 by the merger. Our highest-resolution Hi 21cm images, at a spatial resolution of pc at , find no evidence of Hi at the location of FRB 20250316A or in the neighbouring regions that show star-formation in recent KCWI images. While the location of FRB 20250316A is away from the location of the south-western starburst, the merger event is likely to have triggered star-formation activity across the galaxy (resulting in the higher estimate of the recent SFR), possibly causing the formation of the massive stellar progenitor of FRB 20250316A. We obtain a upper limit of erg s-1 Hz-1 on the rest-frame 1.4 GHz specific luminosity of a persistent radio source associated with FRB 20250316A, one of the strongest current constraints on the radio luminosity of a PRS associated with a non-repeating fast radio burst.

### v2

@@ III Results and Discussion
Applying the above value and the G23 extinction model to our measured H line flux density, we obtain a dust-corrected H line luminosity of erg s-1. Using the calibration of Kennicutt and Evans (2012), this yields an SFR of yr-1 at the location of the eBOSS spectrum, kpc from the centre of the galaxy, for a Chabrier initial mass function. The eBOSS fibres are circular, with diameter , implying an SFR surface density of yr-1 kpc-2 for the region covered by the eBOSS fibre. This is an uncharacteristically high SFR surface density for the outskirts of a disk galaxy, and suggests that the recent burst of star-formation is likely to have been triggered by a merger or accretion event.
Figure 5[A] shows the Galaxy Evolution Explorer (GALEX; Martin et al., 2005) near-ultraviolet (NUV) image of NGC 4141, in colourscale, with the velocity-integrated Hi 21cm emission from the -resolution cube overlaid in contours. Panel [B] of the figure shows a zoom-in on the dashed maroon square shown in [A], without the Hi 21cm contours. While NUV emission is indeed detected around the location of the eBOSS spectrum (indicated by the arrow), this region does not stand out in any way. The measured NUV flux at the above location is erg s-1 cm-2 Å-1, after convolving the image to an angular resolution of , where the errors are from Poisson statistics. For the G23 extinction model and , this yields an SFR of yr-1. This is significantly lower than the SFR of yr-1 from the same location from the H line. The apparent discrepancy can be explained by the fact that the mean age of the stellar population that contributes to the H line emission is Myr, while that contributing to the NUV continuum emission is Myr (e.g. Kennicutt and Evans, 2012). Thus, a very recent episode of star-formation, on timescales Myr would yield a higher SFR from the H line than the estimate from the NUV continuum. The lower SFR estimated from the GALEX NUV image of NGC 4141 is thus consistent with a burst of star-formation, triggered by a recent merger or accretion event. In passing, we note that the angular resolution of the GALEX NUV image is a factor of larger than the diameter of the eBOSS fibres. The SFR inferred from the NUV image is thus averaged over a larger region than that traced by the H emission.
@@
We next used the “direct-Te” method to infer the metallicity at the location of the eBOSS spectrum (e.g. Izotov et al., 2006; Jiang et al., 2019), using the extinction-corrected flux densities in the [Oiii]4363Å, [Oiii]4959Å, [Oiii]5007Å, and [Oii]3727Å lines. This yields . The metallicity at the galaxy centre is (Kewley et al., 2005). This would imply a metallicity gradient of dex kpc-1 towards the south-west. Typical metallicity gradients in disk galaxies are dex kpc-1 (e.g. Zaritsky et al., 1994). A similar gradient, dex kpc-1, is obtained in NGC 4141 from the slightly sub-solar metallicity measured by Chime/FRB Collaboration (2025b) in the FRB environment. Therefore, the low metallicity at the location of the eBOSS spectrum instead suggests a recent addition of metal-poor gas, via either a merger or accretion from the circumgalactic medium. This fresh gas is likely to have resulted in the increased Hi column density seen in the south-western outskirts of the galaxy in Figs. 4[A] and [B], and to have triggered the recent star-formation activity, leading to the high inferred SFR surface density.
@@
Leroy et al. (2019) find that NGC 4141 lies dex above the local star-forming main sequence. Our estimate of the 1.38 GHz flux density ( mJy) allows us to obtain an independent radio-based estimate of the SFR of NGC 4141. Using the Tully-Fisher distance estimate of Mpc (Tully et al., 2023) for the galaxy, we obtain a radio-based SFR of year-1. This is lower than the SFR estimate of year-1 from the dust-corrected H line luminosity (Kewley et al., 2005). As noted by Chime/FRB Collaboration (2025b), this difference may arise due to a recent increase in the star-formation activity, as the radio SFR estimate is an average over 100 Myr, while the H-based estimate is an average over the last 10 Myr. The increased recent SFR may account for the detections of two Type-II SNe, an optical transient, an X-ray transient, and FRB 20250316A, all within the last two decades, as well as for the fact that the galaxy lies above the local main sequence.
@@
Overall, our GMRT Hi 21cm images find evidence for disturbances in the Hi distribution of NGC 4141 and the companion, LEDA 2582852, a nearby Hi cloud to the south-west without any discernible star-formation and at the velocities of the south-western side of the galaxy disk, and high Hi column densities in the south-western outskirts of the galaxy disk of NGC 4141. The SDSS spectrum of the south-western region of the galaxy yields a high SFR surface density and a low metallicity, while the recent ( Myr) SFR of NGC 4141 is higher than the radio-based SFR averaged over 100 Myrs. Further, the SFR inferred from the GALEX NUV image at the location of the eBOSS spectrum is significantly lower than that inferred from the H line. All of these indicate that NGC 4141 has recently acquired fresh gas, via either accretion or a merger, which compressed the Hi in the south-western part of the galaxy, triggering the starburst. This merger/accretion event may have also triggered additional star-formation activity in the disk of NGC 4141, resulting in the increased recent SFR (from the comparison between the radio and the H SFRs of the galaxy) and possibly in the formation of the stellar progenitor of FRB 20250316A, and the progenitors of the other transients. Earlier Hi 21cm mapping studies of FRB host galaxies have also yielded evidence for recent merger activity in a number of systems (e.g Kaur et al., 2022; Glowacki et al., 2024; Roxburgh et al., 2025), although we note that Roxburgh et al. (2025) also found cases where the FRB host galaxy showed no clear evidence for a merger. It is possible that the merger (merger or accretion event, in the case of FRB 20250316A) triggered a burst of star-formation activity that resulted in the formation of the massive stellar progenitors of the ensuing fast radio bursts. We emphasize, however, that the possibility that the progenitor of FRB 20250316A was formed independent of the merger/accretion event cannot be ruled out.
@@
Finally, our non-detection of continuum emission at the location of FRB 20250316A places a upper limit of 24 Jy on the 1.38 GHz flux density of any associated persistent radio source (PRS). This implies the upper limit of erg s-1 on the rest-frame 1.38 GHz radio luminosity of such a putative PRS. This is one of the strongest constraints on the radio luminosity of a PRS associated with a non-repeating FRB (e.g. Law et al., 2024b; Chime/FRB Collaboration, 2025b; An et al., 2025), and approximately four orders of magnitude lower than the radio luminosities of the four repeating FRBs that show a PRS (e.g. Chatterjee et al., 2017; Niu et al., 2022; Bruni et al., 2025; Zhang et al., 2025). For example (for our assumed Tully-Fisher distance of 33.2 Mpc to NGC 4141), An et al. (2025) obtained erg s-1, while Chime/FRB Collaboration (2025b) obtained erg s-1. As noted by An et al. (2025), the stringent constraints on the flux density of a putative PRS associated with FRB 20250316A suggest, for pulsar wind nebula models (e.g. Dai et al., 2017), an evolved progenitor in a low-density, weakly magnetized environment, rather than a young magnetar in a dense environment.
IV Summary
@@
We have used the GMRT Band-5 receivers to map Hi 21cm emission in and around NGC 4141, the host of FRB 20250316A, at , at angular resolutions of , i.e. spatial resolutions of kpc. The low redshift of NGC 4141 implies that this is the finest spatial resolution at which Hi 21cm emission has ever been mapped from an FRB host galaxy. The GMRT Hi 21cm images find evidence for a companion galaxy, LEDA 2582852, to the south-west of NGC 4141, a nearby Hi cloud also towards the south-west, high Hi column densities in the south-western outskirts of the galaxy disk, disturbances in the Hi distribution in both NGC 4141 and LEDA 2582852, and high velocity dispersions, km s-1, away from the central regions of NGC 4141. All of these suggest a recent merger or accretion event that compressed the Hi in the south-western disk, and resulted in a burst of star-formation activity in the galaxy. Consistent with this, an SDSS spectrum in the south-west yields evidence of a low metallicity, suggesting recent addition of metal-poor gas, via either accretion or a merger. The estimate of the recent SFR in NGC 4141, from the H line, is also higher than our estimate of the radio-based SFR, averaged over 100-Myr timescales, while the SDSS spectrum finds a high SFR surface density in the south-western disk. Finally, the SFR estimate from the H line in the south-western disk is significantly larger than the SFR estimate from the NUV continuum, indicating a very recent burst of star-formation, within the last Myr. We conclude that a merger or accretion event within the last Myr is likely to have compressed the Hi in the south-west of the galaxy and triggered the starburst activity. It is unclear whether the merger or accretion event was with the Hi cloud identified kpc away from NGC 4141 or whether this cloud was ejected from NGC 4141 by the merger/accretion event. Our highest-resolution Hi 21cm images, at a spatial resolution of pc at , find no evidence of Hi at the location of FRB 20250316A or in the neighbouring regions that show star-formation in recent KCWI images. While the location of FRB 20250316A is away from the location of the south-western starburst, the merger/accretion event is likely to have triggered star-formation activity across the galaxy (resulting in the higher galaxy-wide SFR estimate from the H line compared to the radio-based SFR estimate), possibly causing the formation of the massive stellar progenitor of FRB 20250316A. We obtain a upper limit of erg s-1 on the rest-frame 1.38 GHz luminosity of a persistent radio source associated with FRB 20250316A, one of the strongest current constraints on the radio luminosity of a PRS associated with a non-repeating fast radio burst.